OpenStudy (christian_10_):

What is the value of y? 6(5 + 3) = 6 • 5 + 6 • y

OpenStudy (anonymous):

y=3

OpenStudy (christian_10_):

Thanks.

OpenStudy (anonymous):

if 6(5+3)=48 then u have to figure out how to get the other side to equal 48. It's 3 because 6*5=30 and 6*3=18 add those and u get 48

OpenStudy (christian_10_):

thanks!

OpenStudy (anonymous):

welcome